The sum of 2 numbers is 24. Seven times the smaller number is the same as 5 times the larger number. Find the smaller number.
Answer by sarah_adam(201)

Let the two numbers be x and y
let the smaller one be x
Given:
x+y = 24
7*x = 5*y
x = (5/7)*y
(5/7)*y + y = 24
12y/7 = 24
12y = 24*7
y = (24*7)/12 = 14
x = 24 - 14 = 10